OpenStudy (anonymous):

PLEASE HELP!! lim x goes to three (x squared - 9 )/(1/3-1/x)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \lim _{x\to 3} \frac{x^2-9}{\frac{1}{3}-\frac{1}{x}} }\)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \lim _{x\to 3} \frac{x^2-9}{\frac{x}{3x}-\frac{3}{3x}}= \lim _{x\to 3} \frac{x^2-9}{\frac{x-3}{3x}} =\lim _{x\to 3} \frac{3x(x^2-9)}{x-3} }\)

OpenStudy (solomonzelman):

Then, apply the difference of squares to \(x^2-9\), and that would cancel out the \(x-3\). After this do a direct substitution (in other words, just plug in x=3).

OpenStudy (anonymous):

thank you!! one more question

OpenStudy (solomonzelman):

Sure

OpenStudy (anonymous):

lim as x goes to 4 [(square root of x plus 3)-3] / (x-4)

OpenStudy (solomonzelman):

\(\color{#000000}{ \displaystyle \lim _{x \to 4}\frac{\sqrt{x+3}-3}{x-4} }\)

OpenStudy (solomonzelman):

this?

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

Ok, there isn't a way to cancel or simplify this.... I really believe that this limit would just diverge to -∞ and +∞, as you approach from the left and right sides respectively.

OpenStudy (anonymous):

THANK YOU!!!

OpenStudy (solomonzelman):

yw